Q8. G) A child rolls a ball on a level floor 3.5 m to another child. If the ball makes 12.0 revolutions, what is its diameter? (ii) A grinding wheel 0.35 m in diameter rotates at 2200 rpm. Calculate its angular velocity in rad s (in)What are (a) the linear speed and (b) the acceleration of a point on the edge of the grinding F, Volu wheel? 09. 2.998 x A child in a boat throws a 5.30-kg package out horizontally with a speed of 10 ms Calculate the velocity of the boat immediately after, assuming it was initially at rest. The mass of the child is 24.0 kg and the mass of the boat is 35.0 kg. 010 What is the change in entropy of 320 g of steam at 100 C when it is condensed to water at 100°C? T period, tim 1.0 kg of water is heated from 0°C to 100 C. Estimate the change in entropy of the water. ny2 7- 011 (i) A novice skier, starting from rest, slides down an icy frictionless 8.0° incline whose vertical height is 105 m. How fast is she going when she reaches the bottom? (i) Jane, looking for Tarzan, is running at top speed (50 msl) and grabs a vine hanging vertically from a tall tree in the jungle. How high can she swing upward? Does the length of the vine affect your answer?
Answer #1

8.

n times * circumference = linear distance

12* ( pi* d) = 3.5

d = 0.093 m

======

ii)

r = 0.35/2 = 0.175 m

w = ( 2 pi / 60)* 2200

w = 230.27 rad/s

iii)

Linear speed = r w = 0.175* 230.27 = 40.3 m/s

Acceleration, a = v^2/r = 40.3^2 / 0.175 = 9280.5 m/s^2

========

Plz post other problem seperately

Do comment in case any doubt will reply for sure . Goodluck